OpenStudy (dancing_stars23):

how to find similar triangles for this one problem? WSR is similar to TVR. Find X going to draw it

OpenStudy (dancing_stars23):

 WS= 10 SR= is not shown, so that's fine WR=x+2 RT=x+6 TV=20 VR= is not show

OpenStudy (dancing_stars23):

@dtan5457 @aaronq

OpenStudy (dancing_stars23):

@amorfide

OpenStudy (amorfide):

okay so they are similar triangles one is a scale factor of the other you are given length WS=10 length VT=20 you do a ratio of the second shape to the first shape to get the scale factor 20:10 2:1 so the scale factor is 2 therefore you know that 2WR=RT

OpenStudy (amorfide):

2(x+2)=x+6

OpenStudy (dancing_stars23):

oh duh, I feel so stupid right now lol -.- so x=2 right?

OpenStudy (dancing_stars23):

@amorfide

OpenStudy (amorfide):

yes! :D

OpenStudy (dancing_stars23):

thank you! ^.^